OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to y=-5/6x+2 and passes through (12, -4).

OpenStudy (aroub):

ah! i hate perpendicular!!! and does the perpindicular lines has to be = to -1 if yes i can help

OpenStudy (anonymous):

^^ aroub idk, i think so!

OpenStudy (anonymous):

almost aroub - if the slope of a line is m then the slope of perpendicular = -1/m

OpenStudy (aroub):

ya it has to be = to -1

OpenStudy (anonymous):

so its y=-1 ??

OpenStudy (aroub):

idk i really forgot the perpendicular

OpenStudy (anonymous):

no the slope of the perpendicular linr wiil be -1/(-5/6) which simplifies to 6/5

OpenStudy (aroub):

ya i think your right because if u simplified it it will be = -1

OpenStudy (anonymous):

so its y=-1x+-5/6 ?

OpenStudy (anonymous):

no the 5/6 is the coefficient of x i.e. (6/5)x the equation u want is y = (6/5)x + 2

OpenStudy (anonymous):

your AWSOME (: (: thanks alot (:

OpenStudy (anonymous):

the general form of the slope-intercept equation is y = mx +c where m = slope of the line and c is the intercept on the y-axis

OpenStudy (anonymous):

no probs bailey

OpenStudy (anonymous):

thanks (: can you help with one last problem like this?

OpenStudy (anonymous):

yes ok

OpenStudy (anonymous):

Write the equation of a line in slope-intercept form that is perpendicular to 3x+2y=6 and passes through the point (3, 0).

OpenStudy (anonymous):

thanks (:

OpenStudy (anonymous):

first manipulate the equation to get it in y=mx+c form divide by 2 (because we have 2x): 3x/2 + y = 3 y = -(3/2)x + 3 -so we can see that the slope = -(3/2) and the intercept on the y axis 3. slope of perpndclr. = -1/-(3/2) = 2/3 so our equation is y = (2/3)x + 3

OpenStudy (anonymous):

sorry on line 2 I should have put 2y not 2x

OpenStudy (anonymous):

to get rid of the fraction we can multiply each term of the equation b 3 3y = 2x + 9 this is equivalent to the fractional equation

OpenStudy (anonymous):

thankss (:

OpenStudy (anonymous):

the questions are a bit strange though - you do not need the coordinates of the point to work out the problem

OpenStudy (anonymous):

okay (:

OpenStudy (anonymous):

if you have a similar problem which gives you the equation and a point on the line you use another form of the formula i e y - y1 = m(x - x1) where m = slope of the line and (x1,y1) are the coordinates of the given point

OpenStudy (aroub):

yes this is the easiest way to do it i agree:)